OpenStudy (jiteshmeghwal9):

Let t be the trace of the matrix \[\Large A=\left[\begin{matrix}\frac{|x+y|}{|x|+|y|} & \alpha_1 & \beta_1 \\ \alpha_2 & \frac{|y+z|}{|y|+|z|}& \beta_2 \\ \alpha_3 & \beta_3 & \frac{|z+x|}{|z|+|x|}\end{matrix}\right]\] then 1). \(0 \le t \le 3 \) 2). \(1 \le t \le 2\) 3). \(1 \le t \le 3\) 4). \(-1 \le t \le 1\)

OpenStudy (jiteshmeghwal9):

@ganeshie8

OpenStudy (inkyvoyd):

the trace is the sum of the diagonals

OpenStudy (inkyvoyd):

also, consider https://en.wikipedia.org/wiki/Triangle_inequality

OpenStudy (reemii):

1) obviously the trace is positive 2) (@inkyvoyd) Take \(x=y=z=1\), what is the trace? 3) is the lower bound 0 or 1? \(\underline{\text{Hint}}\): if \(a\) and \(b\) have the same sign, then \(\frac{|a+b|}{|a|+|b|} = 1\).

OpenStudy (eliesaab):

The right answer is I believe \[ 1\leq t\le3 \]

OpenStudy (eliesaab):

If x,y,z have the some sign then t=3 if two of them are positive say x and y, then \[\frac{\left| x+y\right| }{\left| x\right| +\left| y\right| }+\frac{\left| x+z\right| }{\left| x\right| +\left| z\right| }+\frac{\left| y+z\right| }{\left| y\right| +\left| z\right| }\geq \frac{\left| x+z\right| }{\left| x\right| +\left| z\right| }+\frac{\left| y+z\right| }{\left| y\right| +\left| z\right| }+1\geq 1 \] By symmetry this will do it

OpenStudy (reemii):

That's also how I would finish it @eliesaab @jiteshmeghwal9

OpenStudy (jiteshmeghwal9):